# AnesSuite: A Comprehensive Benchmark and Dataset Suite for Anesthesiology Reasoning in LLMs

### v3

@@ Abstract
The application of large language models (LLMs) in the medical field has garnered significant attention, yet their reasoning capabilities in more specialized domains like anesthesiology remain underexplored. To bridge this gap, we introduce AnesSuite, the first comprehensive dataset suite specifically designed for anesthesiology reasoning in LLMs. The suite features AnesBench, an evaluation benchmark tailored to assess anesthesiology-related reasoning across three levels: factual retrieval (System 1), hybrid reasoning (System 1.x), and complex decision-making (System 2). Alongside this benchmark, the suite includes three training datasets that provide an infrastructure for continued pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning with verifiable rewards (RLVR). Leveraging this suite, we develop Morpheus, the first baseline model collection for anesthesiology reasoning. Despite undergoing limited training with SFT and group relative policy optimization (GRPO), Morpheus demonstrates substantial performance improvements, rivaling the performance of larger-scale models. Furthermore, through comprehensive evaluations and experiments, we analyze the key factors influencing anesthesiology reasoning performance, including model characteristics, training strategies and training data. Both AnesSuite and Morpheus will be open-sourced at https://github.com/MiliLab/AnesSuite.
@@ 4.1 Evaluation
Tab. 4.1 presents the main evaluation results on the AnesBench. We observe that complex anesthesiology reasoning remains a significant challenge for contemporary LLMs. On the AnesBench, most open-source models achieve scores below 0.5 on System 2 questions. Even proprietary models, such as GPT-4o and Claude-3.7-Sonnet, only achieve accuracies of around 0.5 to 0.6. Furthermore, a consistent trend across most models is that performance on System 1.x and System 2 tasks is markedly lower than on System 1 tasks. This discrepancy suggests that the principal challenge for current LLMs is not a simple deficit in anesthesiological knowledge, but rather an underdeveloped ability to apply this knowledge to complex reasoning problems. Notably, medical specific LLMs do not demonstrate a significant advantage over recent general-purpose reasoning models in our evaluation. This finding underscores the specialized nature of anesthesiological reasoning, distinguishing it from general medical domain.
4.2 Baseline Model
The evaluation results for Morpheus are presented in Tab. 4. A key observation is that, after GRPO training, the model achieves consistent improvements across various systems, demonstrating the effectiveness of the GRPO algorithm in enhancing reasoning capabilities. Notably, following SFT and GRPO training on AnesR1, Morpheus-7B, Morpheus-14B and Morpheus-32B reach performance levels comparable to Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct and Qwen2.5-72B-Instruct, respectively. This highlights the efficacy of AnesR1 in cultivating complex reasoning within the specialized domain of anesthesiology. Furthermore, as shown in Tab 17 (Appendix G.1), Morpheus exhibits enhanced reasoning capabilities across a wide range of general medical and general-domain benchmarks (e.g., MMLU, MedQA). Given that AnesR1 focuses exclusively on the niche field of anesthesiology, these results may offer new insights into the the field of reasoning generalization (Cheng et al., 2025).
@@ C.2 Experiment Details for Baseline Model
Furthermore, we trained multiple models using the AnesR1 dataset with the number of training steps extended to 200. As shown in Tab 11, the results demonstrate significant performance improvements across both the Chinese and English subsets of AnesBench, validating the high quality and rich knowledge encapsulated within AnesR1. Notably, as shown in Tab. 4, the Qwen model exhibited a decline in Chinese subset following SFT, which may be attributed to its inherent sensitivities.
@@ Appendix E Case Study
Case 2: Cardiogenic Shock Question: A 70-year-old man presents to the hospital with chest pain and dyspnea. ECG shows 2-$m$ ST elevations in V2-V4. His vital signs are as follows: HR 85, BP 100/55, SpO2 95% on room air. Laboratory test results are remarkable for troponin 3.5 ng/mL and lactate 6 mmol/L. Emergent cardiac catheterization is planned. What other treatment should be initiated at this time? A. Dopamine B. 1 L crystalloid bolus C. Dobutamine D. Norepinephrine E. No other intervention necessary besides immediate cardiac catheterization Correct Answer: C Gemini-2.5-Pro (Incorrect): This is a step-by-step analysis of the question and options. 1. **Analyze the Patient’s Clinical Presentation:** * **Diagnosis:** The patient presents with classic symptoms of an acute myocardial infarction (chest pain, dyspnea). The ECG confirms an ST-segment elevation myocardial infarction (STEMI) in the anterior leads (V2-V4), which typically involves the left anterior descending (LAD) artery. The elevated troponin level further confirms myocardial necrosis. * **Hemodynamics:** The patient is hypotensive with a blood pressure of 100/55 mmHg. The mean arterial pressure (MAP) is approximately 70 mmHg [(2*55 + 100)/3]. While a MAP of 70 might seem borderline acceptable, the systolic pressure is low, and more importantly, there is evidence of end-organ hypoperfusion. * **Evidence of Shock:** The elevated lactate level of 6 mmol/L is a critical finding. It indicates significant tissue hypoperfusion and anaerobic metabolism, which is the hallmark of shock. * **Synthesis:** The combination of a large anterior STEMI (a common cause of pump failure), hypotension, and elevated lactate establishes the diagnosis of cardiogenic shock. The primary problem is the heart’s inability to pump effectively, leading to inadequate cardiac output and systemic hypoperfusion. 2. **Evaluate the Treatment Options:** * **A. Dopamine:** Dopamine is a vasopressor and inotrope. However, clinical trials (e.g., the SOAP II trial) have shown that in patients with cardiogenic shock, dopamine is associated with a higher incidence of arrhythmias and potentially higher mortality compared to norepinephrine. Therefore, it is no longer considered the first-line agent for cardiogenic shock. * **B. 1 L crystalloid bolus:** In cardiogenic shock, the primary problem is pump failure, not volume depletion. The left ventricle is already failing and likely has a high filling pressure (left ventricular end-diastolic pressure). Administering a large fluid bolus would further increase this pressure, potentially worsening pulmonary edema (as suggested by the patient’s dyspnea) without improving cardiac output, according to the Frank-Starling mechanism in a failing heart. Fluid administration should be done with extreme caution, if at all, in cardiogenic shock. * **C. Dobutamine:** Dobutamine is a primary inotrope (increases contractility) with some vasodilatory effects (due to beta-2 agonism). While it can improve cardiac output, its vasodilatory properties can worsen hypotension, which is dangerous in a patient who is already hypotensive. It is often used after blood pressure has been stabilized with a vasopressor or in patients with cardiogenic shock who are not significantly hypotensive. Starting dobutamine as the initial agent in this patient would be inappropriate. * **D. Norepinephrine:** Norepinephrine is the first-line vasopressor recommended by current guidelines (e.g., European Society of Cardiology, American Heart Association) for the management of cardiogenic shock. It is a potent alpha-1 agonist, which causes peripheral vasoconstriction and increases systemic vascular resistance (SVR), thereby raising blood pressure. It also has modest beta-1 agonist effects, which can provide some inotropic support without causing the significant tachycardia or arrhythmogenicity associated with dopamine. By increasing diastolic blood pressure, it also improves coronary perfusion pressure, which is crucial in the setting of an MI. Initiating a norepinephrine infusion is the most appropriate step to stabilize the patient’s blood pressure and improve organ perfusion while awaiting definitive revascularization. * **E. No other intervention necessary besides immediate cardiac catheterization:** While emergent cardiac catheterization is the definitive treatment to address the underlying cause of the shock (the occluded coronary artery), the patient is hemodynamically unstable. Delaying hemodynamic support until arrival in the catheterization lab is unsafe and increases the risk of cardiac arrest and further end-organ damage. The patient must be stabilized pharmacologically as a bridge to revascularization. **Conclusion:** The patient is in cardiogenic shock secondary to an anterior STEMI. The immediate priority, alongside preparing for emergent revascularization, is to correct the hypotension and improve tissue perfusion. Norepinephrine is the recommended first-line vasopressor for this purpose. **Final Answer:** The final answer is **D** Morpheus-14B (Correct): First, the patient is a 70-year-old man with chest pain and dyspnea. The ECG shows ST elevations in V2-V4, which suggests an anterior STEMI (ST-elevation myocardial infarction). His vital signs are: HR 85 (which is normal), BP 100/55 (hypotensive, as diastolic is low), SpO2 95 He’s scheduled for emergent cardiac catheterization, which is appropriate for STEMI. Now, the question is: what other treatment should be initiated at this time? The options are: A. Dopamine B. 1 L crystalloid bolus C. Dobutamine D. Norepinephrine E. No other intervention necessary besides immediate cardiac catheterization I need to think about the patient’s condition. He has hypotension (BP 100/55), which is concerning. Hypotension in a STEMI patient can indicate cardiogenic shock or low perfusion. Lactate is elevated at 6 mmol/L, which is high. Normal lactate is less than 2 mmol/L, so this suggests significant tissue hypoperfusion, likely due to the heart attack. In STEMI, especially with hypotension and elevated lactate, we need to address the shock state. The goal is to improve perfusion and support the heart. Let’s evaluate each option: A. Dopamine: Dopamine is a vasopressor that can increase heart rate and contractility at higher doses. It’s used in shock states, but it’s not first-line for cardiogenic shock from STEMI. Dopamine can have arrhythmogenic effects and isn’t preferred in acute coronary syndromes. B. 1 L crystalloid bolus: This is fluid resuscitation. In hypotension, fluids can be given to increase preload and cardiac output. However, in cardiogenic shock from STEMI, fluid boluses are generally avoided because they can worsen pulmonary edema or increase left ventricular end-diastolic pressure. The patient has dyspnea, which might indicate heart failure or pulmonary issues, so fluids could be harmful. C. Dobutamine: Dobutamine is a beta-agonist inotropic agent that increases cardiac contractility and heart rate. It’s often used in cardiogenic shock to support cardiac output. In STEMI with hypotension, dobutamine can be beneficial to improve perfusion while waiting for reperfusion. D. Norepinephrine: Norepinephrine is a potent vasopressor that increases systemic vascular resistance and blood pressure. It’s commonly used in septic shock or other distributive shocks, but in cardiogenic shock, it can be used if other agents fail. However, it might not be the first choice because it can cause vasoconstriction, which might increase afterload and reduce coronary perfusion. E. No other intervention necessary: But the patient is hypotensive and has elevated lactate, indicating shock. Immediate catheterization is planned, but we need to stabilize him before or during transport. The patient is going for emergent catheterization, so interventions should support him until reperfusion. In STEMI management guidelines, for patients with cardiogenic shock, inotropes like dobutamine are recommended to support hemodynamics while PCI is performed. Also, lactate level of 6 mmol/L is high, suggesting severe shock. We need to address this urgently. Option B, crystalloid bolus, is not ideal because it could worsen the situation. Option A, dopamine, is less preferred due to potential side effects. Option D, norepinephrine, might be used if dobutamine isn’t sufficient, but it’s not first-line for cardiogenic shock from STEMI. The most appropriate choice is C, dobutamine, because it directly supports cardiac contractility without increasing afterload excessively. In some protocols, norepinephrine is used, but dobutamine is more specific for inotropy in heart failure. The patient has dyspnea, which could be due to pulmonary edema, so inotropes might help by increasing cardiac output and reducing congestion. Also, hypotension needs correction. So, I think C. Dobutamine is the best choice. But let’s confirm the BP: 100/55. Systolic is 100, which is low for a 70-year-old, and diastolic is 55, which is very low. This indicates poor perfusion. In STEMI with hypotension, dobutamine is often initiated to support the heart. Moreover, lactate is high, so we need to improve perfusion. Therefore, the answer should be C. Dobutamine. The patient presents with ST-elevation myocardial infarction (STEMI) confirmed by ECG, hypotension (BP 100/55), elevated troponin (3.5 ng/mL), and significantly elevated lactate (6 mmol/L), indicating cardiogenic shock. The emergent cardiac catheterization is appropriate, but additional interventions are needed to support hemodynamics and improve tissue perfusion before reperfusion. - **Option A (Dopamine)**: Dopamine is a vasopressor that can increase heart rate and contractility at higher doses but is not first-line for cardiogenic shock in STEMI due to potential arrhythmogenic effects and lack of strong evidence for benefit in this setting. - **Option B (1 L crystalloid bolus)**: Fluid resuscitation is generally avoided in cardiogenic shock from STEMI because it can worsen pulmonary edema or increase left ventricular end-diastolic pressure, exacerbating dyspnea and heart failure. - **Option C (Dobutamine)**: Dobutamine is a beta-agonist inotropic agent that increases cardiac contractility and heart rate, improving cardiac output and perfusion without significantly increasing afterload. It is recommended in guidelines for cardiogenic shock in STEMI to support hemodynamics while awaiting reperfusion. - **Option D (Norepinephrine)**: Norepinephrine is a potent vasopressor that increases systemic vascular resistance and blood pressure but can reduce coronary perfusion by increasing afterload. It is typically reserved for refractory shock or when inotropes like dobutamine are insufficient. - **Option E (No other intervention necessary)**: The patient’s hypotension and elevated lactate indicate shock, requiring immediate intervention to prevent organ damage. Delaying treatment could worsen outcomes. The most suitable choice is **C. Dobutamine**, as it directly addresses the cardiogenic shock by enhancing cardiac contractility, supporting perfusion, and reducing the risk of complications during transport to catheterization. Qwen-2.5-14B-Instruct (Incorrect): Let’s analyze the patient’s condition step by step: 1. **Clinical Presentation**: The patient is a 70-year-old male with chest pain and dyspnea, which are classic symptoms of a possible acute coronary syndrome (ACS), specifically a STEMI (ST-elevation myocardial infarction) given the ECG findings of 2-$m$ ST elevations in V2-V4. 2. **Vital Signs**: HR 85, BP 100/55, SpO2 95 3. **Laboratory Results**: Troponin 3.5 ng/mL (elevated, indicating myocardial damage) and lactate 6 mmol/L (elevated, indicating tissue hypoperfusion and possible shock). 4. **Management**: Given the patient’s presentation, emergent cardiac catheterization is appropriate. However, the patient’s hypotension and elevated lactate suggest the need for immediate hemodynamic support. Now, let’s evaluate each option: A. **Dopamine**: Dopamine can be used in shock states, but it has a dose-dependent effect. At low doses, it can increase renal and splanchnic blood flow, but at higher doses, it can increase myocardial oxygen demand and potentially worsen ischemia. It is not the first-line agent for hypotension in this setting. B. **1 L Crystalloid Bolus**: This is a reasonable initial step in hypotensive patients to assess for volume responsiveness. However, given the patient’s low blood pressure and elevated lactate, it is likely that the patient is in cardiogenic shock rather than hypovolemic shock, making this less likely to be the primary intervention. C. **Dobutamine**: This is a beta-1 selective inotrope that increases cardiac output by increasing myocardial contractility. It is often used in cardiogenic shock to improve cardiac output and perfusion. However, it does not address the hypotension directly. D. **Norepinephrine**: This is a potent alpha-1 agonist that increases systemic vascular resistance and can be used to treat hypotension in cardiogenic shock. It is the most appropriate choice for immediate hemodynamic support in this patient. E. **No other intervention necessary besides immediate cardiac catheterization**: Given the patient’s hypotension and elevated lactate, immediate hemodynamic support is necessary to stabilize the patient before cardiac catheterization. Considering the patient’s hemodynamic instability and the need for immediate support, the most appropriate initial treatment is: **D. Norepinephrine** This choice addresses the patient’s hypotension and provides the necessary hemodynamic support while preparing for emergent cardiac catheterization.

### v4

@@ Abstract
The application of large language models (LLMs) in the medical field has garnered significant attention, yet their reasoning capabilities in more specialized domains like anesthesiology remain underexplored. To bridge this gap, we introduce AnesSuite, the first comprehensive dataset suite specifically designed for anesthesiology reasoning in LLMs. The suite features AnesBench, an evaluation benchmark tailored to assess anesthesiology-related reasoning across three levels: factual retrieval (System 1), hybrid reasoning (System 1.x), and complex decision-making (System 2). Alongside this benchmark, the suite includes three training datasets that provide an infrastructure for continued pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning with verifiable rewards (RLVR). Leveraging this suite, we develop Morpheus, the first baseline model collection for anesthesiology reasoning. Despite undergoing limited training with SFT and group relative policy optimization (GRPO), Morpheus not only achieves substantial improvements in anesthesiology that rival larger-scale models, but also demonstrates enhanced reasoning capabilities across general medical and broad-domain benchmarks. Furthermore, through comprehensive evaluations and experiments, we analyze the key factors influencing anesthesiology reasoning performance, including model characteristics, training strategies and training data. Both AnesSuite and Morpheus will be open-sourced at https://github.com/MiliLab/AnesSuite.
@@ 4.1 Evaluation
Tab. 4.1 presents the main evaluation results on the AnesBench. We observe that complex anesthesiology reasoning remains a significant challenge for contemporary LLMs. On the AnesBench, most open-source models achieve scores below 0.5 on System 2 questions. Even proprietary models, such as GPT-4o and Claude-3.7-Sonnet, only achieve accuracies of around 0.5 to 0.6. Furthermore, a consistent trend across most models is that performance on System 1.x and System 2 tasks is markedly lower than on System 1 tasks. This discrepancy suggests that the principal challenge for current LLMs is not a simple deficit in anesthesiological knowledge, but rather an underdeveloped ability to apply this knowledge to complex reasoning problems. A major factor contributing to these lower scores in complex reasoning tasks is the prevalence of logical hallucinations during the step-by-step generation process. To better understand this phenomenon, we provide a detailed evaluation of reasoning hallucinations, such as Non-Sequiturs and Over-extrapolations, in Appendix G.3. Additionally, medical specific LLMs do not demonstrate a significant advantage over recent general-purpose reasoning models in our evaluation. This finding underscores the specialized nature of anesthesiological reasoning, distinguishing it from general medical domain.
@@
4.2 Baseline Model
The evaluation results for Morpheus are presented in Tab. 4. A key observation is that, after GRPO training, the model achieves consistent improvements across various systems, demonstrating the effectiveness of the GRPO algorithm in enhancing reasoning capabilities. Notably, following SFT and GRPO training on AnesR1, Morpheus-7B, Morpheus-14B and Morpheus-32B reach performance levels comparable to Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct and Qwen2.5-72B-Instruct, respectively. This highlights the efficacy of AnesR1 in cultivating complex reasoning within the specialized domain of anesthesiology. Furthermore, as shown in Tab 18 (Appendix G.1), Morpheus exhibits enhanced reasoning capabilities across a wide range of general medical and general-domain benchmarks (e.g., MMLU, MedQA). However, on specific subsets of these benchmarks, the improvements exhibited by Morpheus are less pronounced. This may be associated with the correlation between the generalization of reasoning capabilities acquired through reinforcement learning and the extent of data exposure during the pre-training phase, as discussed by (Cheng et al., 2025). Furthermore, considering that AnesR1 focuses exclusively on anesthesiology, a domain typically underrepresented in standard LLM training corpora, these findings may offer valuable insights into another pertinent question: whether reasoning skills acquired via reinforcement learning in sparsely exposed domains can effectively transfer to more general domains.
@@ 7 Conclusion
Acknowledgements
This work was supported in part by the New Generation Artificial Intelligence-National Science and Technology Major Project (Grant No. 2025ZD0123602), the National Natural Science Foundation of China (Grant No. U23A20318 and 62276195), the Science and Technology Major Project of Hubei Province (Grant No. 2024BAB046 and 2025BCB026), the Foundation for Innovative Research Groups of Hubei Province (Grant No. 2024AFA017), the Medical Health Science and Technology Project of Zhejiang Province (Grant No.WKJ-ZJ-26004) and the Major Science and Technology Special Plan Project, Department of Science and Technology of Yunnan Province (Grant No. 202502AS080002). This work was also supported by WHU-Kingsoft Joint Lab. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
We would like to thank Professor Bo Du for his helpful discussions throughout this work. Furthermore, we extend our gratitude to Tao Huang and Zhiwei Wang for their assistance in the data collection process.
@@ C.2 Experiment Details for Baseline Model
Furthermore, we trained multiple models using the AnesR1 dataset with the number of training steps extended to 200. As shown in Tab 11, the results demonstrate significant performance improvements across both the Chinese and English subsets of AnesBench, validating the high quality and rich knowledge encapsulated within AnesR1. Notably, as shown in Tab. 4, the Qwen model exhibited a decline in the Chinese subset following SFT, which may be attributed to an alignment tax resulting from an insufficient number of SFT steps intended for domain alignment.
@@
Finally, we conducted supplementary experiments using the InternLM3-8B-Instruct (Cai et al., 2024) model. The training dataset employed was AnesR1, consistent with the setup used to train the Morpheus model in the main text. The number of training steps for SFT and GRPO was set to 100 and 60, respectively. As illustrated in Tab. 12, following SFT and GRPO training on AnesR1, the model demonstrated significant improvements across all three cognitive dimensions, thereby indicating the strong adaptability of our dataset across diverse foundation models.
Model System1 System1.x System2 Overall
InternLM3-8B-Instruct 0.59 0.45 0.39 0.54
InternLM3-8B-Instruct (SFT + RL) 0.62 0.50 0.43 0.58
Table 12: Results of fine-tuned InternLM3 on AnesBench-English
@@ Appendix E Case Study
Case 2: Cardiogenic Shock Question: A 70-year-old man presents to the hospital with chest pain and dyspnea. ECG shows 2-$m$ ST elevations in V2-V4. His vital signs are as follows: HR 85, BP 100/55, SpO2 95% on room air. Laboratory test results are remarkable for troponin 3.5 ng/mL and lactate 6 mmol/L. Emergent cardiac catheterization is planned. What other treatment should be initiated at this time? A. Dopamine B. 1 L crystalloid bolus C. Dobutamine D. Norepinephrine E. No other intervention necessary besides immediate cardiac catheterization Correct Answer: C Gemini-2.5-Pro (Incorrect): This is a step-by-step analysis of the question and options. 1. **Analyze the Patient’s Clinical Presentation:** * **Diagnosis:** The patient presents with classic symptoms of an acute myocardial infarction (chest pain, dyspnea). The ECG confirms an ST-segment elevation myocardial infarction (STEMI) in the anterior leads (V2-V4), which typically involves the left anterior descending (LAD) artery. The elevated troponin level further confirms myocardial necrosis. * **Hemodynamics:** The patient is hypotensive with a blood pressure of 100/55 mmHg. The mean arterial pressure (MAP) is approximately 70 mmHg [(2*55 + 100)/3]. While a MAP of 70 might seem borderline acceptable, the systolic pressure is low, and more importantly, there is evidence of end-organ hypoperfusion. * **Evidence of Shock:** The elevated lactate level of 6 mmol/L is a critical finding. It indicates significant tissue hypoperfusion and anaerobic metabolism, which is the hallmark of shock. * **Synthesis:** The combination of a large anterior STEMI (a common cause of pump failure), hypotension, and elevated lactate establishes the diagnosis of cardiogenic shock. The primary problem is the heart’s inability to pump effectively, leading to inadequate cardiac output and systemic hypoperfusion. 2. **Evaluate the Treatment Options:** * **A. Dopamine:** Dopamine is a vasopressor and inotrope. However, clinical trials (e.g., the SOAP II trial) have shown that in patients with cardiogenic shock, dopamine is associated with a higher incidence of arrhythmias and potentially higher mortality compared to norepinephrine. Therefore, it is no longer considered the first-line agent for cardiogenic shock. * **B. 1 L crystalloid bolus:** In cardiogenic shock, the primary problem is pump failure, not volume depletion. The left ventricle is already failing and likely has a high filling pressure (left ventricular end-diastolic pressure). Administering a large fluid bolus would further increase this pressure, potentially worsening pulmonary edema (as suggested by the patient’s dyspnea) without improving cardiac output, according to the Frank-Starling mechanism in a failing heart. Fluid administration should be done with extreme caution, if at all, in cardiogenic shock. * **C. Dobutamine:** Dobutamine is a primary inotrope (increases contractility) with some vasodilatory effects (due to beta-2 agonism). While it can improve cardiac output, its vasodilatory properties can worsen hypotension, which is dangerous in a patient who is already hypotensive. It is often used after blood pressure has been stabilized with a vasopressor or in patients with cardiogenic shock who are not significantly hypotensive. Starting dobutamine as the initial agent in this patient would be inappropriate. * **D. Norepinephrine:** Norepinephrine is the first-line vasopressor recommended by current guidelines (e.g., European Society of Cardiology, American Heart Association) for the management of cardiogenic shock. It is a potent alpha-1 agonist, which causes peripheral vasoconstriction and increases systemic vascular resistance (SVR), thereby raising blood pressure. It also has modest beta-1 agonist effects, which can provide some inotropic support without causing the significant tachycardia or arrhythmogenicity associated with dopamine. By increasing diastolic blood pressure, it also improves coronary perfusion pressure, which is crucial in the setting of an MI. Initiating a norepinephrine infusion is the most appropriate step to stabilize the patient’s blood pressure and improve organ perfusion while awaiting definitive revascularization. * **E. No other intervention necessary besides immediate cardiac catheterization:** While emergent cardiac catheterization is the definitive treatment to address the underlying cause of the shock (the occluded coronary artery), the patient is hemodynamically unstable. Delaying hemodynamic support until arrival in the catheterization lab is unsafe and increases the risk of cardiac arrest and further end-organ damage. The patient must be stabilized pharmacologically as a bridge to revascularization. **Conclusion:** The patient is in cardiogenic shock secondary to an anterior STEMI. The immediate priority, alongside preparing for emergent revascularization, is to correct the hypotension and improve tissue perfusion. Norepinephrine is the recommended first-line vasopressor for this purpose. **Final Answer:** The final answer is **D** Morpheus-14B (Correct): First, the patient is a 70-year-old man with chest pain and dyspnea. The ECG shows ST elevations in V2-V4, which suggests an anterior STEMI (ST-elevation myocardial infarction). His vital signs are: HR 85 (which is normal), BP 100/55 (hypotensive, as diastolic is low), SpO2 95% on room air (good oxygenation). Lab results: troponin 3.5 ng/mL (elevated, indicating myocardial injury), lactate 6 mmol/L (high, indicating tissue hypoperfusion or shock). He’s scheduled for emergent cardiac catheterization, which is appropriate for STEMI. Now, the question is: what other treatment should be initiated at this time? The options are: A. Dopamine B. 1 L crystalloid bolus C. Dobutamine D. Norepinephrine E. No other intervention necessary besides immediate cardiac catheterization I need to think about the patient’s condition. He has hypotension (BP 100/55), which is concerning. Hypotension in a STEMI patient can indicate cardiogenic shock or low perfusion. Lactate is elevated at 6 mmol/L, which is high. Normal lactate is less than 2 mmol/L, so this suggests significant tissue hypoperfusion, likely due to the heart attack. In STEMI, especially with hypotension and elevated lactate, we need to address the shock state. The goal is to improve perfusion and support the heart. Let’s evaluate each option: A. Dopamine: Dopamine is a vasopressor that can increase heart rate and contractility at higher doses. It’s used in shock states, but it’s not first-line for cardiogenic shock from STEMI. Dopamine can have arrhythmogenic effects and isn’t preferred in acute coronary syndromes. B. 1 L crystalloid bolus: This is fluid resuscitation. In hypotension, fluids can be given to increase preload and cardiac output. However, in cardiogenic shock from STEMI, fluid boluses are generally avoided because they can worsen pulmonary edema or increase left ventricular end-diastolic pressure. The patient has dyspnea, which might indicate heart failure or pulmonary issues, so fluids could be harmful. C. Dobutamine: Dobutamine is a beta-agonist inotropic agent that increases cardiac contractility and heart rate. It’s often used in cardiogenic shock to support cardiac output. In STEMI with hypotension, dobutamine can be beneficial to improve perfusion while waiting for reperfusion. D. Norepinephrine: Norepinephrine is a potent vasopressor that increases systemic vascular resistance and blood pressure. It’s commonly used in septic shock or other distributive shocks, but in cardiogenic shock, it can be used if other agents fail. However, it might not be the first choice because it can cause vasoconstriction, which might increase afterload and reduce coronary perfusion. E. No other intervention necessary: But the patient is hypotensive and has elevated lactate, indicating shock. Immediate catheterization is planned, but we need to stabilize him before or during transport. The patient is going for emergent catheterization, so interventions should support him until reperfusion. In STEMI management guidelines, for patients with cardiogenic shock, inotropes like dobutamine are recommended to support hemodynamics while PCI is performed. Also, lactate level of 6 mmol/L is high, suggesting severe shock. We need to address this urgently. Option B, crystalloid bolus, is not ideal because it could worsen the situation. Option A, dopamine, is less preferred due to potential side effects. Option D, norepinephrine, might be used if dobutamine isn’t sufficient, but it’s not first-line for cardiogenic shock from STEMI. The most appropriate choice is C, dobutamine, because it directly supports cardiac contractility without increasing afterload excessively. In some protocols, norepinephrine is used, but dobutamine is more specific for inotropy in heart failure. The patient has dyspnea, which could be due to pulmonary edema, so inotropes might help by increasing cardiac output and reducing congestion. Also, hypotension needs correction. So, I think C. Dobutamine is the best choice. But let’s confirm the BP: 100/55. Systolic is 100, which is low for a 70-year-old, and diastolic is 55, which is very low. This indicates poor perfusion. In STEMI with hypotension, dobutamine is often initiated to support the heart. Moreover, lactate is high, so we need to improve perfusion. Therefore, the answer should be C. Dobutamine. The patient presents with ST-elevation myocardial infarction (STEMI) confirmed by ECG, hypotension (BP 100/55), elevated troponin (3.5 ng/mL), and significantly elevated lactate (6 mmol/L), indicating cardiogenic shock. The emergent cardiac catheterization is appropriate, but additional interventions are needed to support hemodynamics and improve tissue perfusion before reperfusion. - **Option A (Dopamine)**: Dopamine is a vasopressor that can increase heart rate and contractility at higher doses but is not first-line for cardiogenic shock in STEMI due to potential arrhythmogenic effects and lack of strong evidence for benefit in this setting. - **Option B (1 L crystalloid bolus)**: Fluid resuscitation is generally avoided in cardiogenic shock from STEMI because it can worsen pulmonary edema or increase left ventricular end-diastolic pressure, exacerbating dyspnea and heart failure. - **Option C (Dobutamine)**: Dobutamine is a beta-agonist inotropic agent that increases cardiac contractility and heart rate, improving cardiac output and perfusion without significantly increasing afterload. It is recommended in guidelines for cardiogenic shock in STEMI to support hemodynamics while awaiting reperfusion. - **Option D (Norepinephrine)**: Norepinephrine is a potent vasopressor that increases systemic vascular resistance and blood pressure but can reduce coronary perfusion by increasing afterload. It is typically reserved for refractory shock or when inotropes like dobutamine are insufficient. - **Option E (No other intervention necessary)**: The patient’s hypotension and elevated lactate indicate shock, requiring immediate intervention to prevent organ damage. Delaying treatment could worsen outcomes. The most suitable choice is **C. Dobutamine**, as it directly addresses the cardiogenic shock by enhancing cardiac contractility, supporting perfusion, and reducing the risk of complications during transport to catheterization. Qwen-2.5-14B-Instruct (Incorrect): Let’s analyze the patient’s condition step by step: 1. **Clinical Presentation**: The patient is a 70-year-old male with chest pain and dyspnea, which are classic symptoms of a possible acute coronary syndrome (ACS), specifically a STEMI (ST-elevation myocardial infarction) given the ECG findings of 2-$m$ ST elevations in V2-V4. 2. **Vital Signs**: HR 85, BP 100/55, SpO2 95% on room air. The blood pressure is notably low (hypotension), which is concerning and suggests possible cardiogenic shock. 3. **Laboratory Results**: Troponin 3.5 ng/mL (elevated, indicating myocardial damage) and lactate 6 mmol/L (elevated, indicating tissue hypoperfusion and possible shock). 4. **Management**: Given the patient’s presentation, emergent cardiac catheterization is appropriate. However, the patient’s hypotension and elevated lactate suggest the need for immediate hemodynamic support. Now, let’s evaluate each option: A. **Dopamine**: Dopamine can be used in shock states, but it has a dose-dependent effect. At low doses, it can increase renal and splanchnic blood flow, but at higher doses, it can increase myocardial oxygen demand and potentially worsen ischemia. It is not the first-line agent for hypotension in this setting. B. **1 L Crystalloid Bolus**: This is a reasonable initial step in hypotensive patients to assess for volume responsiveness. However, given the patient’s low blood pressure and elevated lactate, it is likely that the patient is in cardiogenic shock rather than hypovolemic shock, making this less likely to be the primary intervention. C. **Dobutamine**: This is a beta-1 selective inotrope that increases cardiac output by increasing myocardial contractility. It is often used in cardiogenic shock to improve cardiac output and perfusion. However, it does not address the hypotension directly. D. **Norepinephrine**: This is a potent alpha-1 agonist that increases systemic vascular resistance and can be used to treat hypotension in cardiogenic shock. It is the most appropriate choice for immediate hemodynamic support in this patient. E. **No other intervention necessary besides immediate cardiac catheterization**: Given the patient’s hypotension and elevated lactate, immediate hemodynamic support is necessary to stabilize the patient before cardiac catheterization. Considering the patient’s hemodynamic instability and the need for immediate support, the most appropriate initial treatment is: **D. Norepinephrine** This choice addresses the patient’s hypotension and provides the necessary hemodynamic support while preparing for emergent cardiac catheterization.